OpenStudy (anonymous):

I just started calculus. There is a question I can't seem to grasp. "f(x) = cos2x". Find f(-pi/4). I come out with "cos(-pi/2)", but the book says the answer is "0". Can someone help me out? Thanks.

OpenStudy (kainui):

Well, can you answer this question: f(x) = x^2 find f(3)

OpenStudy (freckles):

yeah you just need to evaluate that cos(-pi/2) thingy

OpenStudy (anonymous):

Yea, F(3) for F(x) = x^2 is: 9.

OpenStudy (kainui):

Awesome, can you try applying this same concept to the problem you were given? Try your best, I'll help you if you mess up, what's important is you try.

OpenStudy (anonymous):

Yes, I came up with cos(-pi/2). This seems to be in simplest form...

OpenStudy (freckles):

you can simplify that more

OpenStudy (freckles):

cos(-pi/2)=cos(pi/2) you could use the unit circle

OpenStudy (freckles):